OpenStudy (anonymous):

I have two questions that I am greatly in need of assistance. I took these in a pretest, and cannot for the life of me understand them. :/ If you could please help me step by step, I would be greatly appreciated. Will metal too! Here they are, simple I know____ The perimeter of a rectangle is 64 units. Can the length x of the rectangle can be 20 units hen its width y is 11 units? Second_ -4(x+1)-3=-3(x-4) For the second one I started off doing: -4(x-1)-3=-3(x-4) -4x-4 -3=-3x-12 +3x +3x -1x-4-3=-12 +4 +4 -1x-3=-12 +3 +3 -1x=-12 -1x -1x

OpenStudy (anonymous):

For the first one I did use the formula: p=2l+2w

OpenStudy (lynfran):

if the perimeter= 64 units with the width y=11 units then; length=(p/2)width so, x=(p/2)y substituting those values we get x= (64/2)(11) x=?

OpenStudy (anonymous):

x = 352 ?

OpenStudy (anonymous):

Maybe I was doing it wrong when I was doing the formula: p=2l*2w 64=(2*20)*(2*11) 40 * 22 40*22 = 880

OpenStudy (anonymous):

Yeah, I still don't understand how you received 352. I'm sorry. :(

OpenStudy (lynfran):

no x=21

OpenStudy (anonymous):

I don't disregard that your answer is wrong, I just can't comprehend it. Maybe another way of explaining it?

OpenStudy (lynfran):

i made a mistake it shd be p=2(l+w) and we want ''l'' so divide both side by 2 we get p/2=l+w now take away w from both sides wee get (p/2)-w=l

OpenStudy (lynfran):

so substituting into that we get (64/2)-11=length

OpenStudy (anonymous):

Ah. Okay, now I can see it. I was so focused on trying to use the length that was provided, that it was incorrect. Thank you.

OpenStudy (anonymous):

Could you help me on my other question if you have time?

OpenStudy (lynfran):

for the second question -4(x+1)-3=-3(x-4)...just making sure its this ..see u wrote (x-1) when u started it..

OpenStudy (anonymous):

Ops, yeah its (x+1) sorry typo

OpenStudy (anonymous):

The second step would start out like this: -4x+-4-3=-3x-12 I know that you need to get all the variables on one side of the equation, preferable the left side.

OpenStudy (lynfran):

ok so -4(x+1)-3=-3(x-4) we can start by multiplying out parentheses so -4x-4-3=-3x+12 simplify -4x-7=-3x+12 add (3x) and (7) to boh sides we get -4x+3x=12+7 simplify -x=19 then divide by -1 x=-19

OpenStudy (anonymous):

Okay: You would finally have -1x-7=12 then adding the seven to both sides of the equation. Finally you would have -1x=19 divide -1x by both sides and get x = -19 Ah, I see the mistake on my part, but thank you so much for your time.

OpenStudy (anonymous):

Have a nice night, thank you again! @ LynFran I made sure I gave you a metal.

OpenStudy (lynfran):

ok... welcome ...